All Systems Operational

About This Site

Welcome to the Crucial status page. This page keeps you up to date with any scheduled or unscheduled maintenance on our network.

Core Network   ? Operational
Blaze Cloud Platform   Operational
Virtual Server Platform   ? Operational
Web Hosting Services   ? Operational
Reseller Hosting Services   ? Operational
Website   ? Operational
Crucial Control   ? Operational
Help Centre   ? Operational
Phone Support   Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Past Incidents
Sep 27, 2017

No incidents reported today.

Sep 26, 2017

No incidents reported.

Sep 25, 2017
Resolved - This incident has been resolved.
Sep 25, 17:03 AEST
Monitoring - We have identified the cause of the outage and currently investigating this further.

All affected customers have been migrated to other hypervisors and are now up and running.
Sep 25, 15:55 AEST
Investigating - Be advised our internal monitoring has indicated that a VPS Hypervisor is down or unavailable. As a result your VPS service that resides on this node may not be accessible or connectable.

We are actively investigating and further updates will follow in the next 60 minutes.
Sep 25, 14:34 AEST
Resolved - This incident has been resolved.
Sep 25, 15:51 AEST
Monitoring - The server is now up and running, We are running MySQLCheck to ensure database consistency.
Sep 25, 15:27 AEST
Update - The server was migrated off the hypervisor which had issues and is currently booting up.
Sep 25, 14:58 AEST
Identified - Be advised our internal monitoring has indicated that the mentioned Shared Hosting server is currently down or unavailable.

This may be impacting access to your website, email services, ftp services and access to the cPanel control panel.

We are actively investigating and further updates will follow in the next 60 minutes.
Sep 25, 14:57 AEST
Sep 24, 2017

No incidents reported.

Sep 23, 2017

No incidents reported.

Sep 22, 2017

No incidents reported.

Sep 21, 2017

No incidents reported.

Sep 20, 2017

No incidents reported.

Sep 19, 2017
Resolved - No further issues have been encountered.

This incident is now resolved.
Sep 19, 01:20 AEST
Identified - All Blaze VPS customer servers have been rebooted and are now online.

We continue to investigate.
Sep 19, 00:08 AEST
Update - The hypervisor has been rebooted. Once online we will begin to power on customer servers and investigate the root cause.
Sep 18, 23:55 AEST
Investigating - We are currently investigating an outage affecting up to 15 Blaze VPS customers.

We will provide updates within the next 15 minutes.
Sep 18, 23:43 AEST
Resolved - Consistency checks have finished successfully and no problems were found.

This incident is now resolved.
Sep 19, 01:14 AEST
Monitoring - We are currently running consistency checks over all databases on both servers.
Sep 19, 00:32 AEST
Update - The affected Reseller servers has been restarted and are now online.

We continue to investigate.
Sep 19, 00:08 AEST
Update - The hypervisor which hosts the affected Reseller servers has been restarted.

We will provide further updates within the next 15 minutes.
Sep 18, 23:57 AEST
Investigating - We are currently investigating an outage affecting Reseller servers rs19-eqx3-syd.ha-node.net and rs26-eqx3-syd.ha-node.net

We will provide updates within the next 15 minutes.
Sep 18, 23:44 AEST
Sep 17, 2017

No incidents reported.

Sep 16, 2017

No incidents reported.

Sep 15, 2017
Completed - The scheduled maintenance has been completed.
Sep 15, 23:08 AEST
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 15, 23:00 AEST
Scheduled - Tonight the Crucial Website and Ticket System will have a short outage while maintenance is performed.

This should not last longer then 5-10 minutes.
Sep 15, 15:23 AEST
Postmortem - Read details
Sep 15, 14:26 AEST
Completed - Emergency maintenance was completed at 4:49AM.
Sep 15, 12:52 AEST
Update - All customer virtual machines are online. A full report of tonight's work is being compiled and will be posted shortly.
Sep 15, 04:49 AEST
Update - Customer virtual machines are being restarted. Further information will be available once all customers are back online.
Sep 15, 04:24 AEST
Update - Customer virtual machines are being shutdown now.
Sep 15, 02:00 AEST
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 14, 22:00 AEST
Scheduled - This evening we will be conducting an emergency maintenance on the SSD VPS hypervisor xnsrv006.oa.apdcsy1.crucial.com.au

We have detected a faulty DRAC module in this server and to maintain the operational reliability, we will be replacing it tonight between 10pm and 4am.

During this time the virtual machines on the hypervisors will be shutdown and the hypervisor will be turned off to replace the faulty DRAC module.

We estimate your service may be offline/inaccessible for up to 1 hour during the maintenance window.

We apologize for the lack of advanced notice and expect this maintenance will ensure stability and reliability of the platform.
Sep 14, 15:31 AEST
Sep 13, 2017

No incidents reported.